OpenStudy (anonymous):

We toss two dice 1,000 times. How many times do we expect to have the sum of the two dice equal to 4? A. about 250 B. about 167 C. about 83 D. about 42

OpenStudy (anonymous):

I would say about 167 times!

OpenStudy (anonymous):

In total there is 36 different ways a pair of dice can be rolled. We must figure that there are three different roles that add up to four, (1,3),(2,2)(3,1). We can figure that for each roll the probability of the sum of the dice will be four is 3/36 = 1/12. You can than figure that for every 12 rolls your sum is going to be four. If we have a thousand rolls, than the probability of rolling a pair that's sum is four is 1000/12 = 83.